OpenStudy (anonymous):

Solve. x2 – 2x = 3. a. x = 4 and x = -1 b. x = -3 and x = 1 c. x = 4 and x = -3 d. x = 3 and x = -1

OpenStudy (anonymous):

x^2-2x-3=0 (x-3)(x+1) x-3=0 and x+1=0 x=3 and x=-1